OpenStudy (anonymous):

Please help!!! I have to do factoring on several equations but I don't know how with only the variable. Example: (x + 14)(x - 3) = 0 Now, I know the answers are -14, and 3, but I need to show the work, and that's where I'm stuck. Please help!

OpenStudy (anonymous):

@HelpBlahBlahBlah @SolomonZelman

OpenStudy (misty1212):

hey!!

OpenStudy (misty1212):

they love "show work" even when it is totally obvious

OpenStudy (anonymous):

Yeah, sucks huh? :/

OpenStudy (misty1212):

\[(x+14)(x-3)=0\\ x+14=0\iff x=-14\\ x-3=0\iff x=3\]

OpenStudy (misty1212):

i.e. set each factor equal to zero, solve for \(x\)

OpenStudy (anonymous):

So you just separate them? It's that easy? Oh my gosh thank you!!!